# Non-Clifford and parallelizable fault-tolerant logical gates on constant and almost-constant rate homological quantum LDPC codes via higher symmetries

### v2

@@ VI.2 Generic scheme for universal logical gate set
In order to apply logical CCZ only on three target logical qubits, we should avoid directly applying traversal T which leads to a collective logical CCZ gate acting on logical qubits. We can use a second ancilla LDPC color code copy (as a specific example the quasi-hyperbolic color code) and setting all the logical qubits in the ancilla copy at state |$\ket{\overline{0}}$, as illustrated in Fig. 24(a). One can then consider the 2 copies of LDPC color codes as 6 copies of LDPC toric codes. We start with three targeted logical qubits labeled as L1, L2 and L3 as shown in Fig. 24(a). The logical qubit L1 is supported on $\Sigma_{g}$ in toric-code copy 2. The logical qubits L2 and L3 are supported on $a^{(i)}\times c$ and $b^{(i)}\times c$ in toric-code copy 1 respectively. We can use the disentangling and re-merging processes introduced in Sec. VI.1 to apply two steps of parallel logical SWAPs (via the combination of logical CZ and logical Hadmard) to shuffle L1 to toric-code copy 3 and then to copy 6, and shuffle L2 (L3) to copy 3 (copy 4) and then to copy 4 (copy 5), as illustrated in Fig. 24(a).
@@ VI.3 Enhancing parallelizability of quasi-hyperbolic codes with thickened Dehn twists
Figure 25: (a) The $13g-$ 1 Dehn twist generators for the mapping class group on a genus-$g$ surface $\Sigma_{g}$. (b,c) The action of the thickened Dehn twist along the 2-cycle $b^{(2)}\times c$ on $\Sigma_{g}\times S^{1}$. (d) The action of the Dehn twist along $a^{(3)}$ on $\Sigma_{g}$.
@@
Figure 27: Strategy to perform logical CNOTs between neighboring tori with label $i$ and $i+1$ respectively using thickened Dehn twists. (a) One first perform two steps of logical SWAPs to shuffle the targeted logical qubits $(a^{(i)}\times c;1)$ and $(b^{(i+1)}\times c;1)$ (for even $i$) to the ancilla color code (toric-code copy 4). (b) In the ancilla color code copy, the logical qubits are now encoded sparsely in an alternating pattern, with the corresponding cycles (1-cycles which are extended to 2-cycles by taking a product with the cycle $c$) being highlighted. All the remaining logical qubits are set in the anicilla state |$\ket{\overline{0}}$.
@@
Now we investigate what logical gates these thickened Dehn twists correspond to. In Fig. 25(a), we show the standard $13g-$ 1 Dehn twists which form a generating set of the mapping class group on genus-$g$ surface MCG($\Sigma_{g}$) [72]. These Dehn twists are acted three types of loops: the first two types are $a^{(i)}$ (red) and $b^{(i)}$ (blue) ($i=1,2,\cdots,g$) introduced before, while the third type are $g-1$ loops connecting neighboring genera, denoted by $f^{(i)}$ (yellow) ($i=1,2,\cdots,g-1$). Note that when consider the mapping classes, all these loops are elements of the 1st-homotopy group instead of 1st homology group (Abelianization of the 1st-homotopy group), and hence have directions which are indicated by arrows.
@@
and acts trivially on any other loops and membranes along the 1st- and 2nd-homology basis of $\Sigma_{g}\times S^{1}$ introduced in Sec. V.1, i.e., $B_{1}=\{a^{(i)},b^{(i)},c|\ i=1,2,\dots,g\}$ and $B_{2}=\{b^{(i)}\times c,a^{(i)}\times c,\Sigma_{g}|\ i=1,2,\dots,g\}$. As illustrated in Fig. 25(c), the thickened Dehn twist does a twist on all the loop $a^{i}$ into $b^{(i)}a^{(i)}$ in the 3-manifold. This thickened Dehn twist induces the following map on the logical operators in toric-code copy $1$ (for simplicity with set the logical qubits in the other 2 copies at |$\ket{\overline{0}}$):
@@
Now we investigate what logical gates the thickened Dehn twists $\tilde{D}_{f^{(i)}\times c}$ correspond to. We first see that all the $g-1$ loops $f^{(i)}$ can be expressed by loops $b^{(j)}$ as
@@
In order to obtain a single independent logical CNOT between the neighboring tori, we can choose a sparse encoding, i.e., set logical qubits labeled by $(a^{(i+1)}\times c;1)$ and $(b^{(i)}\times c;1)$ at state |$\ket{\overline{0}}$ for even $i$ as ancilla logical qubits. In other words, we only store logical information in lobical qubits labeled by $(a^{(i)}\times c;1)$ and $(b^{(i+1)}\times c;1)$ for even $i$, as illustrated in Fig. 27(b). This gives rise to
@@
Now we can use the following strategy to apply parallel logical CNOTs: we can apply logical CNOTs between the two types of tori with the same $i$ label, i.e., $\overline{\text{CNOT}}((a^{(i)}\times c;1),(b^{(i)}\times c;1))$ and $\overline{\text{CNOT}}((b^{(i)}\times c;1),(a^{(i)}\times c;1))$ via $\tilde{D}_{b^{(i)}\times c}$ and $\tilde{D}_{a^{(i)}\times c}$ respectively. Now if we want to apply parallel logical CNOTs between neighboring tori with label $i$ and $i+1$ respectively, we can split them into two groups and perform them in two separate rounds. In order to perform parallel $\overline{\text{CNOT}}((a^{(i)}\times c;1),(b^{(i+1)}\times c;1))$ (for even $i$), we can SWAP all the targeted logical qubits $(a^{(i)}\times c;1)$ and $(b^{(i+1)}\times c;1)$ (for even $i$) into an ancilla LDPC code copy at the same homological cycles with all the other logical qubits in the ancilla code copy set at the |$\ket{\overline{0}}$ state and then perform the sequence of Dehn twists $\tilde{D}_{b^{(i+1)}\times c}\tilde{D}_{b^{(i)}\times c}\tilde{D}_{f^{(i)}\times c}$ on the ancilla LDPC code. as illustrated in as illustrated in Fig. 27(a,b). Afterwards, we SWAP these logical qubits back to the original code copy. Next, in order to perform parallel $\overline{\text{CNOT}}((a^{(i+1)}\times c;1),(b^{(i)}\times c;1))$ (for even $i$), we can SWAP all the targeted logical qubits $(a^{(i+1)}\times c;1)$ and $(b^{(i)}\times c;1)$ into the ancilla LDPC code copy and then perform the sequence of Dehn twists $\tilde{D}_{b^{(i+1)}\times c}\tilde{D}_{b^{(i)}\times c}\tilde{D}_{f^{(i)}\times c}$ on the ancilla LDPC code. The protocol is complete after performing SWAP of logical qubits back to the original code copy.
@@ VI.4 Parallelizability and connectivity of the interaction hypergraph
VII Back-engineering of 3-manifolds from the interaction hypergraph: enhancing parallelizability of logical gates
@@
Consider $\mathsf{H}_{m}$ as $m$ handles attached to the three sphere, with the handles $h_{i}$, $h_{j}$, and $h_{k}$, for $10<i<j<k<m+$ 1, attached to the northern hemisphere and the rest of the $m-3$ handles attached to the southern hemisphere. Let $\sigma_{i,j,k}$ be the mapping class
@@
such that it acts as $\sigma$ from (182) on its restriction to the northern hemisphere and extends by identity to the southern hemisphere. It is clear that $\sigma_{i,j,k}$ belongs to the sub group $G(m)$ hence $H_{2}(\mathsf{H}(\sigma_{i,j,k}),\mathbb{Z})=\mathbb{Z}^{m}$. Moreover, $\mathcal{M}^{3}(\sigma_{i,j,k})$ is a connected sum of $T^{3}$ with $m-3$ copies of $S^{2}\times S^{1}$. In particular, there is only one tuple of three distinct surfaces in $\mathcal{M}^{3}(\sigma_{i,j,k})$, namely the three two cycles of $T^{3}$, such that they have non-trivial intersection and the number of their intersections is exactly one.
@@
[…]
@@
VIII Discussion and outlook
Although we focus on the construction of LDPC color codes supported on 3-manifolds which have transversal $T$ and $S$ gates as the logical gates, one might worry that the color code stabilizers have relatively large weight and the decoding for 3D color code also may be more sophisticated than the 3D toric codes [75]. An alternative approach will be still using 3 copies of homological LDPC codes supported on the same 3-manifolds. Using the disentangling unitary between the single copy of color code and three copies of toric codes in Ref. [24], one can transform the transversal $T$ or $S$ gate acting on the LDPC color codes to a constant-depth local circuit acting on three copies of the homological LDPC code supported on the same 3-manifold which implement the same logical gate. In this way, the homological LDPC code will still have relatively low stabilizer weight corrresponding to the triangulations of the 3-manifold and one can use the usual 3D toric-code decoder or the adapted version for hyperbolic spaces. Another potential way of producing such constant-depth circuits would be to directly translate the cup product expression of the emergent symmetry operator (logical gate) in the discrete state sum Eqs. (43) and (LABEL:eq:CCZ_defect) from the TQFT formalism into gates acted on qubits. We will leave these exploration into future works.
Finally, we point out that although this paper focuses on homological codes defined on manifolds, the triple intersection structure and higher-form symmetries may likely to also exist in quantum expander codes. Therefore, the methods developed in this paper are expected to have a broad application to fault-tolerant quantum computing with qLDPC codes.

### v3

@@ VI.2 Generic scheme for universal logical gate set
[…]
@@ VI.3 Enhancing parallelizability of quasi-hyperbolic codes with thickened Dehn twists
Figure 24: (a) The 3g13g-1 Dehn twist generators for the mapping class group on a genus-$g$ surface $\Sigma_{g}$. (b,c) The action of the thickened Dehn twist along the 2-cycle $b^{(2)}\times c$ on $\Sigma_{g}\times S^{1}$. (d) The action of the Dehn twist along $a^{(3)}$ on $\Sigma_{g}$.
@@
[…]
@@
Now we investigate what logical gates these thickened Dehn twists correspond to. In Fig. 24(a), we show the standard 3g13g-1 Dehn twists which form a generating set of the mapping class group on genus-$g$ surface MCG($\Sigma_{g}$) [75]. These Dehn twists are acted three types of loops: the first two types are $a^{(i)}$ (red) and $b^{(i)}$ (blue) ($i=1,2,\cdots,g$) introduced before, while the third type are g1g-1 loops connecting neighboring genera, denoted by $f^{(i)}$ (yellow) ($i=1,2,\cdots,g-1$). Note that when consider the mapping classes, all these loops are elements of the 1st-homotopy group instead of 1st homology group (Abelianization of the 1st-homotopy group), and hence have directions which are indicated by arrows.
@@
[…]
@@
Now we investigate what logical gates the thickened Dehn twists $\tilde{D}_{f^{(i)}\times c}$ correspond to. We first see that all the g1g-1 loops $f^{(i)}$ can be expressed by loops $b^{(j)}$ as
@@
[…]
@@
We note that the thickened Dehn twists can be performed in parallel, and with all the Dehn twists discussed above we hence can obtain parallel logical CNOTs and SWAPs between all the neighboring logical qubits in the logical quantum circuit. We further note that one can also perform thickened Dehn twists along very non-local membrane beyond those related to the (3gg-1) MCG generator, and in this case logical CNOTs and SWAPs can be performed on logical qubits separated far apart in the logical circuit diagram.
@@ VI.4 Parallelizability and connectivity of the interaction hypergraph
VII Discussion and outlook
Although we focus on the construction of LDPC color codes supported on 3-manifolds which have transversal $T$ and $S$ gates as the logical gates, one might worry that the color code stabilizers have relatively large weight and the decoding for 3D color code also may be more sophisticated than the 3D toric codes [78]. An alternative approach will be still using 3 copies of homological LDPC codes supported on the same 3-manifolds and the corresponding constant-depth circuit corresponding to cup products. In this way, the homological LDPC code will still have relatively low stabilizer weight corrresponding to the triangulations of the 3-manifold and one can use the usual 3D toric-code decoder or the adapted version for hyperbolic spaces.
Finally, we point out that although this paper focuses on homological codes defined on manifolds, the cup-product operation and triple intersection structure, as well as the higher-form symmetries can also be extended to expander-based quantum codes, as shown in the follow-up papers [51, 52]. Therefore, the methods developed in this paper are expected to have a broad application to fault-tolerant quantum computing with qLDPC codes.
Appendix A Back-engineering of 3-manifolds from the interaction hypergraph: enhancing parallelizability of logical gates
@@ VII Discussion and outlook
Consider $\mathsf{H}_{m}$ as $m$ handles attached to the three sphere, with the handles $h_{i}$, $h_{j}$, and $h_{k}$, for $0<i<j<k<m+1$, attached to the northern hemisphere and the rest of the m3m-3 handles attached to the southern hemisphere. Let $\sigma_{i,j,k}$ be the mapping class
@@
such that it acts as $\sigma$ from (173) on its restriction to the northern hemisphere and extends by identity to the southern hemisphere. It is clear that $\sigma_{i,j,k}$ belongs to the sub group $G(m)$ hence $H_{2}(\mathsf{H}(\sigma_{i,j,k}),\mathbb{Z})=\mathbb{Z}^{m}$. Moreover, $\mathcal{M}^{3}(\sigma_{i,j,k})$ is a connected sum of $T^{3}$ with m3m-3 copies of $S^{2}\times S^{1}$. In particular, there is only one tuple of three distinct surfaces in $\mathcal{M}^{3}(\sigma_{i,j,k})$, namely the three two cycles of $T^{3}$, such that they have non-trivial intersection and the number of their intersections is exactly one.
@@
$|\overline{\Gamma}_{i}\cap\overline{\Gamma}_{j}\cap\overline{\Gamma}_{k}|=a_{i,j,k}\,\,mod\,2,$ (178)
@@
Appendix B Proof of Theorem 2
Here, we provide the proof of Theorem 2 in Sec. V.2.
Proof.
The volume of the twisted manifold $\mathcal{M}^{3}_{g}$ scales as
$=$ $vol(\Sigma_{g}\times[0,1])=area(\Sigma_{g})=4\pi(g-1)=O(g),$ (185)
$I(\tau_{*}):=\{v\in H_{1}(\Sigma_{g};\mathbb{Z}_{2})\,|\,\tau_{*}(v)=v\}$ (186)
$\pi:\Sigma_{g}\to\Sigma_{g}/\langle\tau\rangle=:{{}_{g}S}$ (188)
$1\to\pi_{1}(\Sigma_{g})\to\pi_{1}(\Sigma_{g}/\langle\tau\rangle)\xrightarrow{q}\langle\tau\rangle\to 1$ (189)
In particular, the linear combination
$\tilde{\alpha}_{1}+\dots+\tilde{\alpha}_{N(\alpha)}=:v$ (192)
which implies
An easy application of Riemann-Hurwitz formula gives
$=$ $O(g/\log^{\frac{1}{2}}(g))=O(V/\log^{\frac{1}{2}}(V)),$ (196)
where the last equality is obtained via Eq. B.
One can prove the lower bound of the 1-systole $sys_{1}(\mathcal{M}^{3};\mathbb{Z}_{2})$$\Omega(\log^{\frac{1}{2}}(V))$ by contradiction. We assume there exists a non-contractible loop $\gamma$ in $\mathcal{M}^{3}_{g}$ such that $length(\gamma)<O(\log^{\frac{1}{2}}(V))=O(\log^{\frac{1}{2}}(g))$. Consider lifting $\gamma$ in $\mathcal{M}^{3}_{g}$ to a curve $\tilde{\gamma}$ in the covering space $\Sigma_{g}\times\mathbb{R}$. The lifted curve $\tilde{\gamma}$ starts at point $(p,t)$ and ends at $(\tau^{m}p,t+m)$ for some $m\in\mathbb{Z}$. Since the bundle projection $\pi:\mathcal{M}^{3}_{g}\rightarrow[0,1]/\langle 0=1\rangle$ is length non-increasing, we have $|m|\leq length(\tilde{\gamma})=length(\gamma)<O(\log^{\frac{1}{2}}(g))$. Since we have $order(\tau)=\Omega(\log^{\frac{1}{2}}(g))$ from Eq. (89), we obtain $|m|<order(\tau)$, which suggests $p\neq\tau^{m}p$. Note that under the covering map $\sigma:\Sigma_{g}\rightarrow\Sigma_{g}/\langle\tau\rangle=:{{}_{g}S}$, both $p$ and $\tau^{m}p$ in $\Sigma_{g}$ are projected to the same point in $_{g}S$, therefore $p$ and $\tau^{m}p$ must differ by a non-trivial covering translation of the cover $\sigma$. However, any non-trivial covering translation must move each point of the total space $\Sigma_{g}$ by no less than twice the injectivity radius of the base $_{g}S$ since the geodisc balls of radius smaller than the injectivity radius $R(_{g}S)$ must be disjoint. According to Eq. (91), we have $R(_{g}S)\geq O(\log^{\frac{1}{2}}(g_{k}))$. Since the projection $\Sigma_{g}\times\mathbb{R}\rightarrow\Sigma_{g}$ is length nonincreasing, we have
$Area(z)\geq\int_{0}^{1/2}|z\cap\Sigma_{g}\times\{t\}|dt$
So for some $t_{0}\in[0,1/2]$, we have $|z\cap\Sigma_{g}\times\{t_{0}\}|\leq 2Area(z)$. If $z\cap\Sigma_{g}\times\{t_{0}\}$ is non-trivial in $H_{1}(\Sigma_{g};\mathbb{Z}_{2})$ then from $sys_{1}(\Sigma_{g})\gtrsim log(g)$ we get that $Area(z)\gtrsim log(g)$. Otherwise, $z\cap\Sigma_{g}\times\{t_{0}\}$ bounds a 2-cycle $F\subset\Sigma_{g}\times\{t_{0}\}$. Using Buser’s inequality with $\lambda_{1}(\Sigma_{g})\gtrsim 1$, we can find a $F$ with $Area(F)\lesssim Area(z)$. But then the cycle z+FFz+F-F can be lifted to $\Sigma_{g}\times\mathbb{R}$ from which we can see that $Area(z)+2Area(F)\geq Area(\Sigma_{g})\gtrsim g$. From the estimate on $Area(F)$ we get that $Area(z)\gtrsim g$. Thus, we see that $sys_{2}(\mathcal{M}_{g})\gtrsim\log(g)=\Omega(\log V)$. ∎